QD= 1,450 - 25P (Demand)
QS= -100 + 75P (Supply)

Consumer: A: $4.5; B; $3; C: 2; D: 1.25: E: 1
Supplier: S: $3; T: 2.5; U: 2; V: 1.5 W: 1

Calculate the producer surplus under the equilibrium and the price ceiling under $4. What is the deadweight loss for producer and consumer.
1
Expert's answer
2020-10-21T10:18:28-0400

In equilibrium Qd = Qs, so:

1,450 - 25P = -100 + 75P,

P = $15.5,

Q = 1,062.5 units.

If P = $4, then Q = -100 + 75×4 = 200 units.

Producer surplus in equilibrium is:

"PS = 0.5\u00d71,062.5\u00d7(15.5 - 1.33) = 7,527.81."

After the price ceiling is imposed:

"PS = 0.5\u00d7200\u00d7(4 - 1.33) = 267."